OpenStudy (anonymous):

The Probability of wining a particular game at the game place is 1/4. can you find the odds against wining the game. thank you

OpenStudy (anonymous):

probability of wnning +loosing=1

OpenStudy (anonymous):

3/4, or is there more to it?

OpenStudy (anonymous):

yes but not much more

OpenStudy (anonymous):

probability of winning is one fourth , so every four games you should win one and lose three odd in favor are one to three odds against are three to one

OpenStudy (ash2326):

\[\text{odds against}=\frac{\text{Probability of losing}}{\text{Probability of Winning}}=\frac{\frac 3 4}{\frac 14}\]

OpenStudy (anonymous):

to be fancy, odds in favor of event are \(\frac{P(A)}{1-P(A)}\)

OpenStudy (anonymous):

what ash said, although it you think about it the way i wrote, without a formula example: probability of winning is \(\frac{2}{5}\) then odds in favor are two to three and odds against are three to two